OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. just need help setting it up

OpenStudy (anonymous):

When in doubt, draw a picture! Draw the outward journey above the return journey and label all of the forces acting on the plane, then use that to set up your equations.

OpenStudy (anonymous):

S = D/t D = S*t = 40 * 3.5 = 140 m

OpenStudy (anonymous):

Return Trip S = 140/2.5 = 56mph

OpenStudy (anonymous):

@biancataylor515 Did u understand..till here

OpenStudy (anonymous):

this wasn't on the answer sheet though

jimthompson5910 (jim_thompson5910):

Hint: Equation for trip there: d = (r+40)*3.5 Equation for trip back: d = (r-40)*2.5

jimthompson5910 (jim_thompson5910):

where d is the distance traveled and r is the speed of the plane in still air

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

np